# Credal Learning Theory

### v3

@@ 4.1 Realizability and Finite Hypotheses Space
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of credal set $\mathcal{P}$. Let the empirical risk minimizer be
@@
Assume that there exists a realizable hypothesis, that is, $h^{\star}\in\mathcal{H}$ such that $L_{P}(h^{\star})=0$, and that the model space $\mathcal{H}$ is finite. Let $l$ denote the zero-one loss, i.e. $l((x,y),h)=\mathbbm{I}[y\neq h(x)]$. Fix any $\delta\in(0,1)$. Then,
$\mathbb{P}\left[L_{P}(\hat{h})\leq\epsilon^{\star}(\delta)\right]\geq 1-\delta,$
where $\epsilon^{\star}(\delta)$ is a well-defined quantity that depends only on $\delta$ and on extreme elements $\text{ex}\mathcal{P}$ of $\mathcal{P}$, i.e. those that cannot be written as a convex combination of one another.
Under some hypotheses, Theorem 4.1 gives us a tight probabilistic bound for the expected risk $L_{P}(\hat{h})$ of the empirical risk minimizer $\hat{h}$. [^7] Notice how, if $\mathcal{P}$ has finitely many extreme elements – which happens e.g. if we put $\mathcal{P}=\text{Conv}(\{\mathcal{L}_{i}\}_{i=1}^{N})$, another frequentist way of deriving a credal set – then $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}$ is a finite union, and hence it is easier to compute. The bound holds for any possible distribution $P$ in the credal set $\mathcal{P}$ that generates the training data. A slightly looser bound depending on the diameter of credal set $\mathcal{P}$ holds if we calculate $L_{Q}(\hat{h})$ in place of $L_{P}(\hat{h})$.
@@
$\mathbb{P}\left[L_{Q}(\hat{h})\leq\epsilon^{\star}(\delta)+\eta\right]\geq 1-\delta,$
@@
Corollary 4.2 gives us a probabilistic bound for the expected risk $L_{Q}(\hat{h})$ of the empirical risk minimizer $\hat{h}$, calculated with respect to the “wrong” distribution $Q$. That is, any distribution in $\mathcal{P}$ that is different from the one that generates the training set.
We now give a looser – but easier to compute – bound for $L_{P}(\hat{h})$.
@@
Retain the assumptions of Theorem 4.1. We have that
$\epsilon^{\star}(\delta)\leq\epsilon_{\text{UB}}(\delta)\doteq\frac{\log|\mathcal{H}|+\log\left(\frac{1}{\delta}\right)}{n}.$
In turn,
$\mathbb{P}\left[L_{P}(\hat{h})\leq\epsilon_{\text{UB}}(\delta)\right]\geq 1-\delta,$
for all $P\in\Delta_{\mathcal{X}\times\mathcal{Y}}$.
Notice how $\epsilon_{\text{UB}}(\delta)$ is a uniform bound, that is, it holds for all possible distributions on $\mathcal{X}\times\mathcal{Y}$, not just those in $\mathcal{P}$. Strictly speaking, this means that we do not need to come up with a credal set $\mathcal{P}$ to find such a bound. Observe, though, that the bound $\epsilon^{\star}(\delta)$ that we derive in Theorem 4.1 is tighter because it is based on a procedure that leverages the evidence encoded in the credal set $\mathcal{P}$. A synthetic experiment confirming this, and studying other interesting properties of $\epsilon^{\star}(\delta)$ and $\epsilon_{\text{UB}}(\delta)$, can be found in Appendix B. By the proof of Theorem 4.1, we have that $L_{P}(\hat{h})$ behaves as $\mathcal{O}(\log|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}|/n)$, which is faster than the rate $\mathcal{O}(\log|\mathcal{H}|/n)$ that we find in Corollary 4.3, since $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}\subseteq\mathcal{H}$. Roughly, $B_{P^{\text{ex}}}$ is the set of “bad hypotheses” according to $P^{\text{ex}}\in\text{ex}\mathcal{P}$. That is, those $h$’s for which $L_{P^{\text{ex}}}(h)$ is larger than 00. A formal definition is given in the proof of Theorem 4.1. The modeling effort required by producing credal set $\mathcal{P}$ is rewarded with a tighter bound and a faster rate. Notice that Corollary 4.3 corresponds to Liang (2016, Theorem 4): we obtain a classical result as a special case of our more general theorem.
Let us now allow for distribution drift.
@@
$\hat{h}_{1}$ $\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\frac{1}{k}\sum_{i=1}^{k}l((x_{i},y_{i}),h),$
$\hat{h}_{2}$ $\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\frac{1}{n-k}\sum_{i=k+1}^{n}l((x_{i},y_{i}),h).$
@@ 4.2 No Realizability and Finite Hypotheses Space
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of credal set $\mathcal{P}$. Assume that the model space $\mathcal{H}$ is finite. Let $l$ denote the zero-one loss, i.e. $l((x,y),h)=\mathbbm{I}[y\neq h(x)]$. Let also $\hat{h}$ be the empirical risk minimizer, and $h^{\star}$ be the best theoretical model. Fix any $\delta\in(0,1)$. Then,
$\mathbb{P}\left[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)\right]\geq 1-\delta,$
where $\epsilon^{\star\star}(\delta)$ is a well-defined quantity that depends only on $\delta$ and on the elements of $\text{ex}\mathcal{P}$.
We now show that also the “wrong” expected risk $L_{Q}(\hat{h})$ – that is, the expected risk computed according to $Q\in\mathcal{P}$ different from the one generating the training set – concentrates around the expected risk $L_{P}(h^{\star})$ evaluated at the best theoretical model $h^{\star}$.
@@
Retain the assumptions of Theorem 4.5. Denote by $Q\in\mathcal{P}$, $Q\neq P$, a generic distribution in $\mathcal{P}$ different from $P$. Pick any $\eta\in\mathbb{R}_{>0}$; if the diameter $\text{diam}_{TV}(\mathcal{P})=\eta$, we have that
$\mathbb{P}\left[L_{Q}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)+\eta\right]\geq 1-\delta,$
@@
Retain the assumptions of Theorem 4.5. Then,
$\epsilon^{\star\star}(\delta)\leq\epsilon^{\prime}_{\text{UB}}(\delta)\doteq\sqrt{\frac{2\left(\log\left|\mathcal{H}\right|+\log\left(\frac{2}{\delta}\right)\right)}{n}}.$
In turn,
$\mathbb{P}\left[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\prime}_{\text{UB}}(\delta)\right]\geq 1-\delta,$
for all $P\in\Delta_{\mathcal{X}\times\mathcal{Y}}$.
The main difference with respect to Theorem 4.1 is that in Theorem 4.5, $L_{P}(\hat{h})-L_{P}(h^{\star})$ behaves as $\mathcal{O}(\sqrt{\log|B^{\prime}_{\text{ex}\mathcal{P}}|/n})$, which is slower than what we had in Theorem 4.1. This is due to the relaxation of the realizability hypothesis. Just like before, though, we have that $\mathcal{O}(\sqrt{\log|B^{\prime}_{\text{ex}\mathcal{P}}|/n})$ is faster than the rate $\mathcal{O}(\sqrt{\log|\mathcal{H}|/n})$ that we find in Corollary 4.7. This because $B^{\prime}_{\text{ex}\mathcal{P}}\subseteq\mathcal{H}$. Roughly, $B^{\prime}_{\text{ex}\mathcal{P}}$ is the set of “bad hypotheses” according to at least one $P^{\text{ex}}\in\text{ex}\mathcal{P}$. That is, those $h$’s for which $|\hat{L}_{P^{\text{ex}}}(h)-L_{P^{\text{ex}}}(h)|$ is larger than 00, for at least one $P^{\text{ex}}$. A formal definition is given in the proof of Theorem 4.5. Notice that Corollary 4.7 corresponds to Liang (2016, Theorem 7): we obtain a classical result as a special case of our more general theorem.
@@
$\mathbb{P}\bigg{[}\bigg{(}$ $L_{P_{1}}(\hat{h}_{1})-L_{P_{1}}\left(h^{\star}_{P_{1}}\right)\bigg{)}+\bigg{(}L_{P_{2}}(\hat{h}_{2})-L_{P_{2}}\left(h^{\star}_{P_{2}}\right)\bigg{)}$
$\leq\epsilon^{\star\star}(\delta)\sqrt{\frac{n}{k(n-k)}}\left(\sqrt{k}+\sqrt{n-k}\right)\bigg{]}\geq 1-\delta,$
@@
Corollary 4.8 tells us that the excess risk is bounded also in the presence of distribution drift. The price we pay for allowing distribution shift is a looser bound. As a result of Corollary 4.7, for a looser but easier to compute bound, we can substitute $\epsilon^{\star\star}(\delta)$ with $\epsilon^{\prime}_{\text{UB}}(\delta)$.
@@ 4.3 No Realizability and Infinite Hypotheses Space
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of credal set $\mathcal{P}$. Let $l$ denote the zero-one loss, i.e. $l((x,y),h)=\mathbbm{I}[y\neq h(x)]$. Let also $\hat{h}$ be the empirical risk minimizer, and $h^{\star}$ be the best theoretical model. Fix any $\delta\in(0,1)$. Then,
@@
for all $P\in\mathcal{P}$. Here,
$\epsilon^{\star\star\star}(\delta)\doteq 4\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})+\sqrt{\frac{2\log(2/\delta)}{n}},$
where
$\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})\doteq\sup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}R_{n,P^{\text{ex}}}(\mathcal{A})$
and
@@
Here, $\sigma_{1},\ldots,\sigma_{n}\sim\text{Unif}(\{-1,1\})$, and $\mathcal{A}\doteq\{(x,y)\mapsto l((x,y),h):h\in\mathcal{H}\}$. [^8] Class $\mathcal{A}$ is the loss class, and it is the composition of the zero-one loss function with each of the hypotheses in $\mathcal{H}$ (Liang, 2016, Page 70).
@@
Notice that $R_{n,P^{\text{ex}}}(\mathcal{A})$ is a slight modification of the classical definition of the Rademacher complexity of class $\mathcal{A}$ [^9] Recall that the Rademacher complexity of $\mathcal{A}$ measures how well the best element of $\mathcal{H}$ fits random noise (coming from the $\sigma_{i}$’s) (Liang, 2016, Page 69), (Balcan and Berlind, 2011). Indeed, the latter is given by
@@
$R_{n}(\mathcal{A})\equiv R_{n,P}(\mathcal{A})\doteq\mathbb{E}_{P}\left[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}l((x_{i},y_{i}),h)\right].$
@@
This because the expectation is taken with respect to the same distribution $P$ according to which the training elements $(x_{1},y_{1}),\ldots,$ $(x_{n},y_{n})$ are drawn. We consider $R_{n,P^{\text{ex}}}(\mathcal{A})$ instead of $R_{n,P}(\mathcal{A})$ because, since $\mathcal{P}$ is a credal set, we have that $P$ can be written as a convex combination of the extreme elements of $\mathcal{P}$, and $\sup_{P\in\mathcal{P}}R_{n,P}(\mathcal{A})=\sup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}R_{n,P^{\text{ex}}}(\mathcal{A})$. If the credal set is finitely generated, that is, if it has finitely many extreme elements, [^10] See footnote 7. then it is easier to compute $\epsilon^{\star\star\star}(\delta)$: we only need to compute a maximum in place of a supremum.
@@
As we show in Corollary 4.10, Theorem 4.9 generalizes Liang (2016, Theorem 9). This latter focuses only on the “true” probability $P^{\star}$ on $\mathcal{X}\times\mathcal{Y}$, while our result holds for all the plausible distributions in credal set $\mathcal{P}$. This grants us to hedge against distribution misspecification. Let us pause here to add a clarification. In real-world applications, we effectively cannot compute $R_{n,P^{\star}}(\mathcal{A})$, since distribution $P^{\star}$ is unknown. While $R_{n,P^{\star}}(\mathcal{A})$ can be approximated via the empirical Rademacher complexity $\hat{R}_{n}(\mathcal{A})$ (Liang, 2016, Equation (219)), whose expected value is indeed $R_{n,P^{\star}}(\mathcal{A})$, doing so has at least two drawbacks:
(1) Especially in the case of low cardinality training set, i.e., if $n\equiv n_{D_{N+1}}$ is not “large enough”, this may lead to a poor approximation of the classical bound (equation (13) in Appendix A);
(2) The collected dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ may well be a realization from the tail of distribution $P^{\star}$. The empirical Rademacher complexity $\hat{R}_{n}(\mathcal{A})$, then, would be a poor approximation of $R_{n,P^{\star}}(\mathcal{A})$.
Instead, while $\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})$ is more conservative, it can be computed explicitly – since we know credal set $\mathcal{P}$ and its extreme elements $\text{ex}\mathcal{P}$ – and it leads to a bound that, although looser, holds for all $P\in\mathcal{P}$.
@@
Retain the assumptions of Theorem 4.9. If $\mathcal{P}$ is the singleton $\{P^{\star}\}$, we retrieve Liang (2016, Theorem 9).
@@
$\mathbb{P}\left[L_{Q}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star\star}(\delta)+\eta\right]\geq 1-\delta,$
@@
Let us now allow for distribution drift.
@@
Consider a natural number $k<n$. Let $(x_{1},y_{1}),\ldots,(x_{k},y_{k})\sim P_{1}$ i.i.d., and $(x_{k+1},y_{k+1}),\ldots,(x_{n},y_{n})\sim P_{2}$ i.i.d., where $P_{1},P_{2}$ are two generic elements of credal set $\mathcal{P}$. Retain the other assumptions of Theorem 4.9, and let
$\epsilon^{\star\star\star}_{\text{shift}}$ $\doteq 4\left[\overline{R}_{k,P^{\text{ex}}}(\mathcal{A})+\overline{R}_{n-k,P^{\text{ex}}}(\mathcal{A})\right]$
$+\sqrt{\frac{2\log(2/\delta)}{n(n-k)}}\left(\sqrt{n-k}+\sqrt{n}\right).$
Then,
$\mathbb{P}\bigg{[}\bigg{(}$ $L_{P_{1}}(\hat{h}_{1})-L_{P_{1}}\left(h^{\star}_{P_{1}}\right)\bigg{)}$
$+\left(L_{P_{2}}(\hat{h}_{2})-L_{P_{2}}\left(h^{\star}_{P_{2}}\right)\right)\leq\epsilon^{\star\star\star}_{\text{shift}}\bigg{]}\geq 1-\delta,$

### v4

@@ 4.1 Realizability and Finite Hypotheses Space
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of the credal set $\mathcal{P}$. Let the empirical risk minimizer be
@@
Assume that there exists a realizable hypothesis, that is, $h^{\star}\in\mathcal{H}$ such that $L_{P}(h^{\star})=0$, and that the model space $\mathcal{H}$ is finite. Let $l$ denote the zero-one loss, and fix any $\delta\in(0,1)$. Then, $\mathbb{P}[L_{P}(\hat{h})\leq\epsilon^{\star}(\delta)]\geq 1-\delta$, where $\epsilon^{\star}(\delta)$ is a well-defined quantity that depends only on $\delta$ and on extreme elements $\text{ex}\mathcal{P}$ of $\mathcal{P}$, i.e., those that cannot be written as a convex combination of one another.
Under the assumptions of finiteness and realizability, Theorem 4.1 gives us a tight probabilistic bound for the expected risk $L_{P}(\hat{h})$ of the empirical risk minimizer $\hat{h}$. The bound holds for any possible distribution $P$ in the credal set $\mathcal{P}$ that generated the stream of training data. A slightly looser bound depending on the diameter of credal set $\mathcal{P}$ holds if we calculate $L_{Q}(\hat{h})$ in place of $L_{P}(\hat{h})$.
@@
$\mathbb{P}[L_{Q}(\hat{h})\leq\epsilon^{\star}(\delta)+\eta]\geq 1-\delta,$
@@
Corollary 4.2 gives us a probabilistic bound for the expected risk $L_{Q}(\hat{h})$ of the empirical risk minimizer $\hat{h}$, calculated with respect to a “wrong” distribution $Q$ – that is, any distribution in $\mathcal{P}$ different from the one generating the new test set of data $D_{N+1}$. We can also give a looser – but easier to compute – bound for $L_{P}(\hat{h})$.
@@
Under the assumptions of Theorem 4.1, $\epsilon^{\star}(\delta)\leq\epsilon_{\text{UB}}(\delta)\doteq 1/n[\log|\mathcal{H}|+\log(\frac{1}{\delta})]$ and
$\mathbb{P}[L_{P}(\hat{h})\leq\epsilon_{\text{UB}}(\delta)]\geq 1-\delta,\quad\forall P\in\Delta_{\mathcal{X}\times\mathcal{Y}}.$
Notice how $\epsilon_{\text{UB}}(\delta)$ is a uniform bound, that is, a bound that holds for all possible distributions on $\mathcal{X}\times\mathcal{Y}$, not just those in $\mathcal{P}$. Strictly speaking, this means that we do not need to come up with a credal set $\mathcal{P}$ to find such a bound. Observe, though, that the bound $\epsilon^{\star}(\delta)$ in Theorem 4.1 is tighter, as it leverages the training evidence encoded in the credal set $\mathcal{P}$. A synthetic experiment confirming this, and studying other interesting properties of $\epsilon^{\star}(\delta)$ and $\epsilon_{\text{UB}}(\delta)$, can be found in Appendix B.
By the proof of Theorem 4.1, we have that $L_{P}(\hat{h})$ behaves as $\mathcal{O}(\log|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}|/n)$, which is faster than the rate $\mathcal{O}(\log|\mathcal{H}|/n)$ that we find in Corollary 4.3, since $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}\subseteq\mathcal{H}$ [^7] Notice how, if $\mathcal{P}$ has finitely many extreme elements – which happens, e.g., if we put $\mathcal{P}=\text{Conv}(\{\mathcal{L}_{i}\}_{i=1}^{N})$, another frequentist way of deriving a credal set – then $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}$ is a finite union, hence easier to compute. Roughly, $B_{P^{\text{ex}}}$ is the set of “bad hypotheses” according to $P^{\text{ex}}\in\text{ex}\mathcal{P}$. That is, those $h$’s for which $L_{P^{\text{ex}}}(h)$ is larger than 00. A formal definition is given in the proof of Theorem 4.1. The modeling effort required by producing credal set $\mathcal{P}$ is therefore rewarded with a tighter bound and a faster rate.
Notice that Corollary 4.3 corresponds to Liang [44, Theorem 4]: we obtain a classical result as a special case of our more general theorem.
Let us now allow for distribution drift in the new test set of data $D_{N+1}$.
@@
$\hat{h}_{1}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\left\{{\frac{1}{k}\sum_{i=1}^{k}l((x_{i},y_{i}),h)}\right\},\quad\hat{h}_{2}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\left\{{\frac{1}{n-k}\sum_{i=k+1}^{n}l((x_{i},y_{i}),h)}\right\}.$
@@ 4.2 No Realizability and Finite Hypotheses Space
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of the credal set $\mathcal{P}$. Assume that the model space $\mathcal{H}$ is finite. Let $l$ be the zero-one loss, $\hat{h}$ the empirical risk minimizer, and $h^{\star}$ the best theoretical model. Fix any $\delta\in(0,1)$. Then, $\mathbb{P}[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)]\geq 1-\delta$, where $\epsilon^{\star\star}(\delta)$ is a well-defined quantity that depends only on $\delta$ and on the elements of $\text{ex}\mathcal{P}$.
As we did in Section 4.1, we can also show that the “wrong” expected risk $L_{Q}(\hat{h})$ – that is, the expected risk computed according to $Q\in\mathcal{P}$ different from the one generating the new evidence $D_{N+1}$ – concentrates around the expected risk $L_{P}(h^{\star})$ evaluated at the best theoretical model $h^{\star}$.
@@
$\mathbb{P}[L_{Q}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)+\eta]\geq 1-\delta,$
@@
Retain the assumptions of Theorem 4.5. Then, $\epsilon^{\star\star}(\delta)\leq\epsilon^{\prime}_{\text{UB}}(\delta)\doteq\sqrt{\frac{2\left(\log\left|\mathcal{H}\right|+\log\left(\frac{2}{\delta}\right)\right)}{n}}.$ In turn, $\mathbb{P}[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\prime}_{\text{UB}}(\delta)]\geq 1-\delta$, for all $P\in\Delta_{\mathcal{X}\times\mathcal{Y}}$.
The main difference with respect to Theorem 4.1 is that in Theorem 4.5, $L_{P}(\hat{h})-L_{P}(h^{\star})$ behaves as $\mathcal{O}(\sqrt{\log|B^{\prime}_{\text{ex}\mathcal{P}}|/n})$, which is slower than what we had in Theorem 4.1. This is due to the relaxation of the realizability hypothesis. Just like before, though, we have that $\mathcal{O}(\sqrt{\log|B^{\prime}_{\text{ex}\mathcal{P}}|/n})$ is faster than the rate $\mathcal{O}(\sqrt{\log|\mathcal{H}|/n})$ that we find in Corollary 4.7. This is because $B^{\prime}_{\text{ex}\mathcal{P}}\subseteq\mathcal{H}$.
Roughly, $B^{\prime}_{\text{ex}\mathcal{P}}$ is the set of “bad hypotheses” according to at least one $P^{\text{ex}}\in\text{ex}\mathcal{P}$. That is, those $h$’s for which $|\hat{L}(h)-L_{P^{\text{ex}}}(h)|$ is larger than 00, for at least one $P^{\text{ex}}$. A formal definition is given in the proof of Theorem 4.5. Notice that Corollary 4.7 corresponds to Liang [44, Theorem 7]: we obtain a classical result as a special case of our more general theorem.
@@
$\mathbb{P}\big{[}(L_{P_{1}}(\hat{h}_{1})-L_{P_{1}}(h^{\star}_{P_{1}}))+(L_{P_{2}}(\hat{h}_{2})-L_{P_{2}}(h^{\star}_{P_{2}}))\leq\epsilon^{\star\star}(\delta)\sqrt{\frac{n}{k(n-k)}}(\sqrt{k}+\sqrt{n-k})\big{]}\geq 1-\delta,$
@@
Corollary 4.8 tells us that the excess risk is also bounded in the presence of distribution drift. The price we pay for allowing distribution shift is a looser bound. As a result of Corollary 4.7, for a looser but easier to compute bound, we can substitute $\epsilon^{\star\star}(\delta)$ with $\epsilon^{\prime}_{\text{UB}}(\delta)$.
@@ 4.3 No Realizability and Infinite Hypotheses Space
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of credal set $\mathcal{P}$. Let $l$ denote the zero-one loss, $\hat{h}$ the empirical risk minimizer and $h^{\star}$ the best theoretical model. Fix any $\delta\in(0,1)$. Then,
@@
for all $P\in\mathcal{P}$. Here, $\epsilon^{\star\star\star}(\delta)\doteq 4\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})+\sqrt{\frac{2\log(2/\delta)}{n}}$, where $\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})\doteq\sup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}R_{n,P^{\text{ex}}}(\mathcal{A})$ and
@@
In (6), $\sigma_{1},\ldots,\sigma_{n}\sim\text{Unif}(\{-1,1\})$, and $\mathcal{A}\doteq\{(x,y)\mapsto l((x,y),h):h\in\mathcal{H}\}$.
@@
$R_{n,P^{\text{ex}}}(\mathcal{A})$ is a slight modification of the classical Rademacher complexity of class $\mathcal{A}$, [^8] Class $\mathcal{A}$ is the loss class, and it is the composition of the zero-one loss function with each of the hypotheses in $\mathcal{H}$ [44, Page 70]. The Rademacher complexity of $\mathcal{A}$ measures how well the best element of $\mathcal{H}$ fits random noise (coming from the $\sigma_{i}$’s) [9], [44, Page 69]. given by
@@
$R_{n}(\mathcal{A})\equiv R_{n,P}(\mathcal{A})\doteq\mathbb{E}_{P}\left[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}l((x_{i},y_{i}),h)\right],$
@@
where the expectation is taken with respect to the same distribution $P$ from which the data points $(x_{1},y_{1}),\ldots,$ $(x_{n},y_{n})$ are drawn. We consider $R_{n,P^{\text{ex}}}(\mathcal{A})$ instead of $R_{n,P}(\mathcal{A})$ because, since $\mathcal{P}$ is a credal set, $P$ can be written as a convex combination of the extreme elements of $\mathcal{P}$, and $\sup_{P\in\mathcal{P}}R_{n,P}(\mathcal{A})=\sup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}R_{n,P^{\text{ex}}}(\mathcal{A})$. If the credal set is finitely generated, that is, if it has finitely many extreme elements (see footnote 7), then it is easier to compute $\epsilon^{\star\star\star}(\delta)$: we only need to compute a maximum in place of a supremum.
@@
As we show in Corollary 4.10, Theorem 4.9 generalizes Liang [44, Theorem 9]. This latter focuses only on the “true” probability $P^{\star}_{N+1}\equiv P^{\text{true}}$ on $\mathcal{X}\times\mathcal{Y}$, while our result holds for all the plausible distributions in credal set $\mathcal{P}$. This grants us to hedge against distribution misspecification.
Let us pause here to add a clarification. In real-world applications, we effectively cannot compute $R_{n,P^{\text{true}}}(\mathcal{A})$, since the distribution $P^{\text{true}}$ is unknown. While $R_{n,P^{\text{true}}}(\mathcal{A})$ can be approximated via the empirical Rademacher complexity $\hat{R}_{n}(\mathcal{A})$ [44, Equation (219)], whose expected value is indeed $R_{n,P^{\text{true}}}(\mathcal{A})$, doing so has at least two drawbacks: (1) When the number of data points $n\equiv n_{D_{N+1}}$ is not “large enough”, this may lead to a poor approximation of the classical bound (Equation (10) in Appendix A); (2) The test set of data $D_{N+1}=\{(x_{i},y_{i})\}_{i=1}^{n}$ may well be a realization from the tail of distribution $P^{\text{true}}\equiv P^{\star}_{N+1}$. The empirical Rademacher complexity $\hat{R}_{n}(\mathcal{A})$, then, would be a poor approximation of $R_{n,P^{\text{true}}}(\mathcal{A})$. In opposition, while $\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})$ is more conservative, it can be computed explicitly – since we know the credal set $\mathcal{P}$ and its extreme elements $\text{ex}\mathcal{P}$ – and it leads to a bound that, although looser, holds for all $P\in\mathcal{P}$.
@@
Retain the assumptions of Theorem 4.9. If $\mathcal{P}$ is the singleton $\{P^{\text{true}}\}$ (i.e., all the training datasets $D_{1},\ldots,D_{N}$ are generated by the same distribution as the new test set $D_{N+1}$), we retrieve Liang [44, Theorem 9].
@@
$\mathbb{P}[L_{Q}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star\star}(\delta)+\eta]\geq 1-\delta,$
@@
Finally, we once again allow for distribution drift.
@@
Consider a natural number $k<n$. Let $(x_{1},y_{1}),\ldots,(x_{k},y_{k})\sim P_{1}$ i.i.d., and $(x_{k+1},y_{k+1}),\ldots,(x_{n},y_{n})\sim P_{2}$ i.i.d., where $P_{1},P_{2}$ are two generic elements of credal set $\mathcal{P}$. Retain the other assumptions of Theorem 4.9, and let $\epsilon^{\star\star\star}_{\text{shift}}\doteq 4[\overline{R}_{k,P^{\text{ex}}}(\mathcal{A})+\overline{R}_{n-k,P^{\text{ex}}}(\mathcal{A})]+\sqrt{\frac{2\log(2/\delta)}{n(n-k)}}(\sqrt{n-k}+\sqrt{n})$. Then,
$\mathbb{P}[(L_{P_{1}}(\hat{h}_{1})-L_{P_{1}}(h^{\star}_{P_{1}}))+(L_{P_{2}}(\hat{h}_{2})-L_{P_{2}}(h^{\star}_{P_{2}}))\leq\epsilon^{\star\star\star}_{\text{shift}}]\geq 1-\delta,$